# Gauge dependence of momentum running in higher-derivative gravity

### v1

@@ Abstract
Recent works [1, 2] have argued that improved one-loop beta-functions capturing the physical momentum dependence of one-loop corrected higher-derivative gravity theories are the most suitable to describe their high-energy behaviour. This work critically tests the validity of this claim. We compute the explicit gauge dependence of the one-loop momentum running of curvature-squared operators in quadratic gravity and conformal gravity using the background field method. We find them to be gauge dependent, and we discuss the implications of this result for the theory and its physical predictivity.
@@ I Introduction
In two previous works by some of the authors [1, 2], it was claimed that the most appropriate one-loop beta functions to describe the high-energy behaviour of higher-derivative theories, and in particular of quadratic gravity and conformal gravity, are not those identified by UV divergences, but those defined by the logarithmic momentum-dependent form factors in the effective action. While these two definitions give identical results in standard theories with $2$-derivative kinetic terms, the same is not true in higher derivative ones [3].
This difference is the result of on-shell infrared divergences that emerge in the large momentum limit, which cannot be treated with the usual machinery introduced for soft and collinear IR divergences [4]. In fact, after a partial fraction decomposition of the propagator, one can show that the same infrared divergences can be seen as sub-subleading logarithmic terms of the UV divergences in a theory with higher-derivative ghosts [5, 6].
@@
$I_{2}(m_{1},m_{2},p)\sim\begin{cases}2\mathcal{O}^{2}\left[\log\left(\frac{p^{2}}{\Lambda^{2}}\right)+\sum_{i}\frac{m_{i}^{2}}{p^{2}}\log\left(\frac{m_{i}^{2}}{p^{2}}\right)+\frac{m_{1}^{2}m_{2}^{2}}{p^{4}}\sum_{i}\log\left(\frac{m_{i}^{2}}{p^{2}}\right)+O\left(\frac{p^{4}}{m_{i}^{4}}\right)\right]\ &\text{if}\ p^{2}\gg m_{i}^{2}\,;\\[4.30554pt] \frac{1}{m_{1}^{2}-m_{2}^{2}}\left[m_{1}^{2}\log\left(\frac{m_{1}^{2}}{\Lambda^{2}}\right)-m_{2}^{2}\log\left(\frac{m_{2}^{2}}{\Lambda^{2}}\right)\right]+O\left(\frac{p}{m_{i}}\right)&\text{if}\ p^{2}\ll m_{i}^{2}\,,\end{cases}$ (5)
@@
In the present work, we compute the gauge dependence of the momentum induced running using the background field formalism [7, 8], both for the beta functions associated with local covariant operators and for those related to the nonlocal partners of the Einstein-Hilbert term [9]. The gauge dependence of momentum-induced beta functions of quadratic gravity has been computed independently in a recent work [6] with a partially alternative approach.
The rest of the paper is structured as follows. In Section II, we briefly present the theories of quadratic gravity and conformal gravity and introduce the background field method, together with our choice for the gauge parameters. In Section III, we present the gauge-dependent beta functions for the operators quadratic in curvature with at most one inverse power of the $\square$ operator. In Section IV, we analyse this gauge dependence and its source within the background field method. In Section V, we summarize the obtained results and discuss future perspectives in the study of quadratic gravity. Throughout this paper, we work with Euclidean signature. In this way, we avoid possible issues with Wick rotation in the presence of higher-derivative poles [10].
@@ II The framework
We start with quadratic gravity [11]. The classical action of quadratic gravity in the Riemann basis is
$\begin{split}S_{\rm qg}[g_{\mu\nu}]=&\int{\rm d}^{4}x\sqrt{g}\,\Bigl{[}-Z_{N}R+\alpha R^{2}+\beta R^{\mu\nu}R_{\mu\nu}+\gamma R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}\Bigr{]}\,,\end{split}$ (7)
@@
$\begin{split}S_{\rm qg}[g_{\mu\nu}]=&\int{\rm d}^{4}x\sqrt{g}\,\Bigl{[}\,-Z_{N}R+\frac{1}{\xi}R^{2}+\frac{1}{2\lambda}C^{\mu\nu\rho\sigma}C_{\mu\nu\rho\sigma}\,\Bigr{]}\,,\end{split}$ (8)
@@
Quadratic gravity is perturbatively renormalizable [11], and that makes it an interesting candidate as a UV completion of general relativity. However, since it has 4-derivative kinetic terms for the gravitonal perturbations, it suffers from Ostrogradsky instability at the classical level [12], and has problems with, alternatively, unbounded energy spectrum, nonunitarity, or negative definite probabilities in its canonically quantized version [13]. Recently, some approaches to solve this problem have been suggested (see, for example, [14, 15, 16, 17, 18]), and, in most cases, the price to be paid is a loss of microcausality at transplanckian scales.
@@
This step has been done in previous works (see, e.g., [19]), thus we will not report the explicit expressions here.
Due to diffeomorphism symmetry, we need to choose a gauge-fixing condition to remove redundant field configurations from the path integral of $h_{\mu\nu}$. Here, we will consider a fourth-derivative gauge-fixing term which makes the scaling of the graviton propagator homogeneous in the deep UV. In practice, we add the following term to the action of quadratic gravity (see, for instance, [20])
@@
The minimal gauge is chosen such that the $4$-derivative gauge-fixed Hessian takes the simple form $\square^{2}$, see, e.g., Ref. [21]. The minimal gauge was adopted in [1] as part of the computation of running couplings in terms of leading-log contributions. In the present paper, we keep the gauge parameters $g_{1}$, $g_{2}$, $g_{3}$, and $g_{4}$ completely arbitrary in order to study the gauge-dependence of the leading-log contributions to the effective action.
@@
Thanks to the background-field method, the effective action as a functional of the background metric $\bar{g}_{\mu\nu}$ (with $h_{\mu\nu}=0$) is a manifestly covariant functional with structure [22, 23]
@@
The effective action is obtained by integrating out $h$, $\bar{c}$, $c$, and $b$ in the path integral, and the one-loop correction is given by [24, 25, 21]
$\begin{split}\Delta\Gamma_{\text{one-loop}}=\frac{1}{2}{\rm tr}\log H-{\rm tr}\log\Delta_{\rm gh}-\frac{1}{2}{\rm tr}\log Y\,,\end{split}$ (17)
@@
$\begin{split}(\Delta_{\rm gh})^{\mu}{}_{\nu}=\delta^{\mu}_{\nu}\,\bar{\Box}+(2g_{2}+1)\,\bar{\nabla}^{\mu}\bar{\nabla}_{\nu}+\bar{R}^{\mu}{}_{\nu}\,,\end{split}$ (18)
@@
$H^{\mu\nu\alpha\beta}=\frac{\delta^{2}(S[\bar{g}+h]+S_{\rm GF}[h,\bar{g}])}{\delta h_{\mu\nu}\delta h_{\rho\sigma}}\Bigg{|}_{h=0}\,.$ (19)
The problem of calculating the UV divergences of these functional traces is an old one, but it can be accomplished using the general results of [26].
The computation of the IR contributions is considerably more challenging. To compute the IR divergent parts, which, as discussed above, occur only in the functional traces of higher derivative operators, we adopt the strategy that has already been described in [1, 2] and mutuated from [27, 28]. We also expand the background metric around flat spacetime as $\bar{g}_{\mu\nu}=\delta_{\mu\nu}+f_{\mu\nu}$, so that we can compute the background’s 2-point function
$\langle f^{\mu\nu}f^{\rho\sigma}\rangle=\frac{\delta^{2}\Gamma[\bar{g}=\delta+f]}{\delta f_{\mu\nu}\delta f_{\alpha\beta}}\Bigg{|}_{f=0}\,.$ (20)
@@
With the structure of the gauge sector described above, the contributions of the FP ghosts and the LN field are functional traces of second-order operators, so they have only UV divergences, which can be computed following the procedures described, for example, in [26, 24]. We will denote them as $\textbf{\text{UVB}}_{\text{gh}}$ and $\textbf{\text{UVB}}_{Y}$, respectively.
$p$$p$$q+p$$q$
$p$$p$$q$
Figure 1: Diagrams contributing to the 2-point function: bubbles (left) and tadpoles (right). The thin line can be the $h$ propagator or one of the ghosts, the thick line is the $f$ propagator, with momentum $p$.
As for the fluctuations $h_{\mu\nu}$, the story is more complicated. The $1$-loop contribution to the 2-point function can be separated into tadpole diagrams and bubble diagrams depicted in Figure 1. In higher derivative (HD) theories, these diagrams can be both UV and IR divergent. The sum of UV divergences is used to define the “traditional” beta functions, which we have referred to as $\mu$-running in previous works [3, 1, 2], where $\mu$ is, for example, the scale introduced by minimal subtraction. They are identified by the $\log\mu$ terms in dimensional regularization in theories where the IR part of loop integrals is regularized by a mass (e.g., the EH term in quadratic gravity). In this scenario, the IR log-divergences are instead identified by $\log m$ factors. In the background 2-point function, there is only one external momentum $p^{\mu}$ due to overall momentum conservation, so the only possible logarithm containing momenta in the high-energy limit ($p^{2}\gg m^{2}$) is $\log p^{2}$. We extract the momentum running, or $p$-running, from the $\log p^{2}$ terms.
@@
[…]
@@ III Gauge dependence
In this section, we display the dependence of the different beta functions on the gauge parameters. In particular, we will show that the beta function encoding the running with respect to the momentum $p$, as defined in [1, 2], turns out to be explicitly gauge dependent.
@@ III.1 Quadratic Gravity
This expression is gauge-independent and in agreement with the literature [24, 29, 21]. The beta functions encoding the $\mu$-running of $\lambda$ (and $\xi$) are defined as $2\lambda^{2}$ (and $\xi^{2}$) times the coefficient of $C^{2}$ (and $R^{2}$) in the last expression, namely
@@
Here, the beta functions encoding the $p$-running of $\lambda$ (and $\xi$) are defined as $-2\lambda^{2}$ (and $-\xi^{2}$) times the coefficient of $C^{2}$ (and $R^{2}$) in expression (23), namely
@@
$p\frac{\partial\xi(p)}{\partial p}\coloneq\beta^{(p)}_{\lambda}=-\frac{\mathcal{C}_{R^{2}}}{10368\pi^{2}(g_{2}+1)^{2}}\,.$ (26b)
@@
$-\Big{[}\textbf{\text{UVB}}+\textbf{\text{IRB}}+\textbf{\text{UVB}}_{\text{gh}}+\textbf{\text{UVB}}_{Y}\Big{]}_{Z_{N}=0}^{\text{min.\leavevmode\nobreak\ gauge}}$ $=\frac{5\left(\xi^{2}-36\lambda\xi-2520\lambda^{2}\right)}{5760\pi^{2}\xi^{2}}\,R\log(-\square)R$ (28)
@@
which is in accordance with the results of Ref. [1].
Once stated the gauge dependence of the momentum running of the local operators in quadratic gravity, we would like to see whether it is possible to define a gauge invariant running for the masses of the spin-2 ghost and the spin-0 mode, which are, respectively, $m_{2}^{2}=\lambda Z_{N}$ and $m_{0}^{2}=-\xi Z_{N}/6$ in the classical theory. It is well known that the beta function encoding the $\mu$-running of the Einstein-Hilbert term is gauge dependent [30, 31], hence a gauge-independent beta function for the mass parameters is missing. We want to understand whether such a beta function can be defined within the momentum running approach.
@@
In the minimal gauge, we recover the well-known result, see, e.g., Ref. [24]
$\big{[}\textbf{\text{UVB}}+\textbf{\text{Tad}}\big{]}^{\text{min.\leavevmode\nobreak\ gauge}}_{Z_{N}}=-\frac{Z_{N}(30\lambda-\xi)(4\lambda+\xi)}{384\pi^{2}\xi}\,R\,.$ (32)
@@
$\mathcal{C}_{R^{2}}^{Z_{N}}$ = $=\,6\,g_{1}\big{[}2160(g_{2}+1)^{2}\lambda^{2}(116g_{2}(g_{2}+2)(g_{3}-g_{4}+1)+116g_{3}-116g_{4}+127)$ (34)
$+7680(g_{2}+1)^{2}\lambda\xi(8g_{2}(g_{2}+2)(g_{3}-g_{4}+1)+8g_{3}-8g_{4}+9)$
$-\xi^{2}(4g_{2}(g_{2}(40g_{2}(38g_{2}+125)(g_{3}-g_{4}+1)+7(885g_{3}-885g_{4}+913))$
$+3470g_{3}-3470g_{4}+3796)+5(604g_{3}-604g_{4}+717))\big{]}$
$+(g_{3}-g_{4}+1)\big{[}3945600(g_{2}+1)^{4}\lambda^{3}+240(104g_{2}(10g_{2}+17)+1025)(g_{2}+1)^{2}\lambda^{2}\xi$
$+40(8g_{2}(218g_{2}+127)+55)(g_{2}+1)^{2}\lambda\xi^{2}-(16g_{2}(g_{2}(8g_{2}(85g_{2}+217)+1671)+773)+2663)\xi^{3}\big{]}\,,$
@@
$\mathcal{C}_{C^{2}}^{Z_{N}}$ $=(-g_{3}+g_{4}-1)\big{[}-51840(g_{2}+1)^{3}(127g_{2}+139)\lambda^{3}-720(g_{2}+1)^{2}(16(g_{2}-34)g_{2}-227)\lambda^{2}\xi$ (35)
$+24(g_{2}+1)(g_{2}(8g_{2}(70g_{2}+243)+1893)+617)\lambda\xi^{2}+(4g_{2}+1)^{2}(64g_{2}(g_{2}+2)+73)\xi^{3}\big{]}$
$-54g_{1}\big{[}-48(g_{2}+1)^{2}\lambda^{2}(1276g_{2}(g_{2}+2)(g_{3}-g_{4}+1)+1276g_{3}-1276g_{4}+1461)$
$+128(g_{2}+1)^{2}\lambda\xi(8g_{2}(g_{2}+2)(g_{3}-g_{4}+1)+8g_{3}-8g_{4}+9)+\xi^{2}(4g_{2}(g_{2}(24g_{2}(2g_{2}+7)(g_{3}-g_{4}+1)$
$+237g_{3}-237g_{4}+241)+6(27g_{3}-27g_{4}+28))+180g_{3}-180g_{4}+191)\big{]}\,.$
@@
$\big{[}\textbf{\text{UVB}}-\textbf{\text{IRB}}\big{]}_{Z_{N}}^{\text{min.\leavevmode\nobreak\ gauge}}$ $=\frac{Z_{N}\left(41868\lambda^{3}+5802\lambda^{2}\xi-178\lambda\xi^{2}-21\xi^{3}\right)}{10368\pi^{2}\xi^{2}}\,R\left(\frac{\log(-\square)}{-\Box}\right)R$ (36)
@@
which is gauge independent, in agreement with Refs. [25, 32]. Thus, the beta function capturing the $\mu$-running of $\lambda$ is given by
@@
$-\Big{[}\textbf{\text{UVB}}+\textbf{\text{IRB}}+\textbf{\text{UVB}}_{\text{gh}}+\textbf{\text{UVB}}_{Y}\Big{]}^{\text{min.\leavevmode\nobreak\ gauge}}=\frac{93}{320\pi^{2}}\,C_{\mu\nu\alpha\beta}\log(-\square)C^{\mu\nu\alpha\beta}\,.$ (42)
@@
$\beta_{\lambda}^{(p)}\big{|}_{\text{min.\leavevmode\nobreak\ gauge}}=-\frac{1}{(4\pi)^{2}}\frac{93\lambda^{2}}{5}\,,$ (43)
which agrees with the result given in Ref. [2].

### v2

@@ Abstract
Recent works [Buccio:2024hys, Buccio:2024omv] have argued that improved one-loop beta-functions capturing the physical momentum dependence of one-loop corrected higher-derivative gravity theories are the most suitable to describe their high-energy behaviour. This work critically tests the validity of this claim. We compute the explicit gauge dependence of the one-loop momentum running of curvature-squared operators in quadratic gravity and conformal gravity using the background field method. We find them to be gauge dependent, and we discuss the implications of this result for the theory and its physical predictivity.
@@ I Introduction
In two previous works by some of the authors [Buccio:2024hys, Buccio:2024omv], it was claimed that the most appropriate one-loop beta functions to describe the high-energy behaviour of higher-derivative theories, and in particular of quadratic gravity and conformal gravity, are not those identified by UV divergences, but those defined by the logarithmic momentum-dependent form factors in the effective action. While these two definitions give identical results in standard theories with $2$-derivative kinetic terms, the same is not true in higher derivative ones [Buccio:2023lzo]. See also [Kawai:2025wkp] for a discussion on the relation between the different definitions of running and possible redefinitions of the scale $\mu$ introduced in dimensional regularization.
This difference is the result of on-shell infrared divergences that emerge in the large momentum limit, which cannot be treated with the usual machinery introduced for soft and collinear IR divergences [Weinberg:1965nx]. In fact, after a partial fraction decomposition of the propagator, one can show that the same infrared divergences can be seen as sub-subleading logarithmic terms of the UV divergences in a theory with higher-derivative ghosts [Buccio:2025his, Salvio:2025cmi].
@@
[…]
@@
In the present work, we compute the gauge dependence of the momentum induced running using the background field formalism [DeWitt:1967ub, Weinberg:1996kr], both for the beta functions associated with local covariant operators and for those related to the nonlocal partners of the Einstein-Hilbert term [Donoghue:2022chi]. The gauge dependence of momentum-induced beta functions of quadratic gravity has been computed independently in a recent work [Salvio:2025cmi] with a partially alternative approach.
The rest of the paper is structured as follows. In Section II, we briefly present the theories of quadratic gravity and conformal gravity and introduce the background field method, together with our choice for the gauge parameters. In Section III, we present the gauge-dependent beta functions for the operators quadratic in curvature with at most one inverse power of the $\square$ operator. In Section IV, we analyse this gauge dependence and its source within the background field method. In Section V, we summarize the obtained results and discuss future perspectives in the study of quadratic gravity. Throughout this paper, we work with Euclidean signature. In this way, we avoid possible issues with Wick rotation in the presence of higher-derivative poles [Aglietti:2016pwz, Buoninfante:2025klm].
@@ II The framework
We start with quadratic gravity [Stelle:1976gc]. The classical action of quadratic gravity in the Riemann basis is
[…]
@@
[…]
@@
Quadratic gravity is perturbatively renormalizable [Stelle:1976gc], and that makes it an interesting candidate as a UV completion of general relativity. However, since the gravitonal perturbations have 4-derivative kinetic terms, the theory suffers from Ostrogradsky unbounded energy at the classical level [Ostrogradsky:1850fid]. Surprisingly, it has been recently shown that an unbounded energy does not always lead to classical instability in higher derivative theories [Deffayet:2023wdg, Held:2025fii]. This feature leads to an unbounded energy spectrum in the canonically quantized version of the theory, and to problems with unitarity, negative probabilities, or recovering classical gravitation when non-canonical quantizations are employed [Woodard:2015zca]. In the past years, some alternative approaches toward a consistent quantization procedure have been suggested (see, for example, [Anselmi:2018ibi, Mannheim:2020ryw, Donoghue:2021eto, Salvio:2015gsi, Holdom:2015kbf]). In most cases, the price to be paid is a loss of microcausality at transplanckian scales.
@@
This step has been done in previous works (see, e.g., [Percacci:2017fkn]), thus we will not report the explicit expressions here.
Due to diffeomorphism symmetry, we need to choose a gauge-fixing condition to remove redundant field configurations from the path integral of $h_{\mu\nu}$. Here, we will consider a fourth-derivative gauge-fixing term which makes the scaling of the graviton propagator homogeneous in the deep UV. In practice, we add the following term to the action of quadratic gravity (see, for instance, [Ohta:2016jvw])
@@
The minimal gauge is chosen such that the $4$-derivative gauge-fixed Hessian takes the simple form $\square^{2}$, see, e.g., Ref. [Ohta:2013uca]. The minimal gauge was adopted in [Buccio:2024hys] as part of the computation of running couplings in terms of leading-log contributions. In the present paper, we keep the gauge parameters $g_{1}$, $g_{2}$, $g_{3}$, and $g_{4}$ completely arbitrary in order to study the gauge-dependence of the leading-log contributions to the effective action.
@@
Thanks to the background-field method, the effective action as a functional of the background metric $\bar{g}_{\mu\nu}$ (with $h_{\mu\nu}=0$) is a manifestly covariant functional with structure [Barvinsky:1990up, Codello:2012kq]
@@
The effective action is obtained by integrating out $h$, $\bar{c}$, $c$, and $b$ in the path integral, and the one-loop correction is given by [Avramidi:1985ki, Fradkin:1981iu, Ohta:2013uca]
[…]
@@
[…]
@@
$H^{\mu\nu\alpha\beta}=\frac{\delta^{2}(S[\bar{g}+h]+S_{\rm GF}[h,\bar{g}])}{\delta h_{\mu\nu}\delta h_{\alpha\beta}}\Bigg|_{h=0}\,.$ (19)
The problem of calculating the UV divergences of these functional traces is an old one, but it can be accomplished using the general results of [Barvinsky:1985an].
The computation of the IR contributions is considerably more challenging. To compute the IR divergent parts, which, as discussed above, occur only in the functional traces of higher derivative operators, we adopt the strategy that has already been described in [Buccio:2024hys, Buccio:2024omv] and mutuated from [tHooft:1974toh, Julve:1978xn]. We also expand the background metric around flat spacetime as $\bar{g}_{\mu\nu}=\delta_{\mu\nu}+f_{\mu\nu}$, so that we can compute the background’s 2-point function
$\langle f^{\mu\nu}f^{\alpha\beta}\rangle=\frac{\delta^{2}\Gamma[\bar{g}=\delta+f]}{\delta f_{\mu\nu}\delta f_{\alpha\beta}}\Bigg|_{f=0}\,.$ (20)
@@
With the structure of the gauge sector described above, the contributions of the FP ghosts and the LN field are functional traces of second-order operators, so they have only UV divergences, which can be computed following the procedures described, for example, in [Barvinsky:1985an, Avramidi:1985ki]. We will denote them as $\textbf{\text{UVB}}_{\text{gh}}$ and $\textbf{\text{UVB}}_{Y}$, respectively.
ppppq+pq+pqq
ppppqq
Figure 1: Diagrams contributing to the 2-point function: bubbles (left) and tadpoles (right). The thin line can be the $h$ propagator or one of the ghosts, the thick line is the $f$ propagator, with momentum pp.
As for the fluctuations $h_{\mu\nu}$, the story is more complicated. The $1$-loop contribution to the 2-point function can be separated into tadpole diagrams and bubble diagrams depicted in Figure 1. In higher derivative (HD) theories, these diagrams can be both UV and IR divergent. The sum of UV divergences is used to define the “traditional” beta functions, which we have referred to as $\mu$-running in previous works [Buccio:2023lzo, Buccio:2024hys, Buccio:2024omv], where $\mu$ is, for example, the scale introduced by minimal subtraction. They are identified by the $\log\mu$ terms in dimensional regularization in theories where the IR part of loop integrals is regularized by a mass (e.g., the EH term in quadratic gravity). In this scenario, the IR log-divergences are instead identified by $\log m$ factors. In the background 2-point function, there is only one external momentum $p^{\mu}$ due to overall momentum conservation, so the only possible logarithm containing momenta in the high-energy limit ($p^{2}\gg m^{2}$) is $\log p^{2}$. We extract the momentum running, or $p$-running, from the $\log p^{2}$ terms.
@@
The bubble integral depends instead on $p$ and can produce $\log p$ terms, so there is no such trivial relation between its UV and IR divergent parts. In the following, we will divide the one-loop contribution to the effective action into three terms: the UV bubble, the IR bubble, and the tadpole. We will call UVB the coefficient in front of $\log\mu$ in the UV bubble, Tad the same coefficient in the tadpole, and IRB the coefficient of $\log m$ in the IR part of the bubble diagram. Given these definitions, in the graviton part of the functional trace, the contribution to the $\mu$-running is given by the sum $\textbf{\text{UVB}}+\textbf{\text{Tad}}$, while the quantity $(\textbf{\text{UVB}}+\textbf{\text{IRB}})$, which precisely corresponds to the coefficient of $\log|p|$, contributes to the $p$-running. Finally, the sum of IR divergences is the difference $\textbf{\text{IRB}}-\textbf{\text{Tad}}$. It is quite easy to check that, in this way, the arguments of all logarithms can be made dimensionless. Furthermore, each of these three linear combinations must be covariant, while the same is not true for UVB, IRB, and Tad taken independently.
@@ III Gauge dependence
In this section, we display the dependence of the different beta functions on the gauge parameters. In particular, we will show that the beta function encoding the running with respect to the momentum $p$, as defined in [Buccio:2024hys, Buccio:2024omv], turns out to be explicitly gauge dependent.
@@ III.1 Quadratic Gravity
This expression is gauge-independent and in agreement with the literature [Avramidi:1985ki, deBerredo-Peixoto:2004cjk, Ohta:2013uca]. The beta functions encoding the $\mu$-running of $\lambda$ (and $\xi$) are defined as $2\lambda^{2}$ (and $\xi^{2}$) times the coefficient of $C^{2}$ (and $R^{2}$) in the last expression, namely
@@
Here, the beta functions encoding the $p$-running of $\lambda$ (and $\xi$) are defined as $2\lambda^{2}$ (and $\xi^{2}$) times the coefficient of $C^{2}$ (and $R^{2}$) in expression (23), namely
@@
$p\frac{\partial\xi(p)}{\partial p}\coloneq\beta^{(p)}_{\xi}=-\frac{\mathcal{C}_{R^{2}}}{10368\pi^{2}(g_{2}+1)^{2}}\,.$ (26b)
@@
$-\Big[\textbf{\text{UVB}}+\textbf{\text{IRB}}+\textbf{\text{UVB}}_{\text{gh}}+\textbf{\text{UVB}}_{Y}\Big]_{Z_{N}=0}^{\text{min.\penalty 10000\ gauge}}$ $=\frac{5\left(\xi^{2}-36\lambda\xi-2520\lambda^{2}\right)}{5760\pi^{2}\xi^{2}}\,R\log(-\square)R$ (28)
@@
which is in accordance with the results of Ref. [Buccio:2024hys].
Once stated the gauge dependence of the momentum running of the local operators in quadratic gravity, we would like to see whether it is possible to define a gauge invariant running for the masses of the spin-2 ghost and the spin-0 mode, which are, respectively, $m_{2}^{2}=\lambda Z_{N}$ and $m_{0}^{2}=-\xi Z_{N}/6$ in the classical theory. It is well known that the beta function encoding the $\mu$-running of the Einstein-Hilbert term is gauge dependent [Avramidi:1986mj, Shapiro:1994vb], hence a gauge-independent beta function for the mass parameters is missing. We want to understand whether such a beta function can be defined within the momentum running approach.
@@
In the minimal gauge, we recover the well-known result, see, e.g., Ref. [Avramidi:1985ki]
$\big[\textbf{\text{UVB}}+\textbf{\text{Tad}}\big]^{\text{min.\penalty 10000\ gauge}}_{Z_{N}}=-\frac{Z_{N}(30\lambda-\xi)(4\lambda+\xi)}{384\pi^{2}\xi}\,R\,.$ (32)
@@
[…]
@@
[…]
@@
$\big[\textbf{\text{UVB}}-\textbf{\text{IRB}}\big]_{Z_{N}}^{\text{min.\penalty 10000\ gauge}}$ $=\frac{Z_{N}\left(41868\lambda^{3}+5802\lambda^{2}\xi-178\lambda\xi^{2}-21\xi^{3}\right)}{10368\pi^{2}\xi^{2}}\,R\left(\frac{\log(-\square)}{-\Box}\right)R$ (36)
@@
which is gauge independent, in agreement with Refs. [Fradkin:1981iu, deBerredo-Peixoto:2003jda, Ohta:2015zwa]. Thus, the beta function capturing the $\mu$-running of $\lambda$ is given by
@@
[…]
@@
$\beta_{\lambda}^{(p)}\big|_{\text{min.\penalty 10000\ gauge}}=-\frac{1}{(4\pi)^{2}}\frac{93\lambda^{2}}{5}\,$, (43)
which agrees with the result given in Ref. [Buccio:2024omv].